Solve the equation.
[1/8]^-2p=2^3p+2
a. p=2/3
b. p=-2/9
c. p=-9/2
d. p=3/2

Answer by chiefman(11)

(1/8)^-2p=2^3p+2
1/8 becomes 2^(-3) since 1/a=a^(-1)
hence the eqn becomes 2^(-3)*-2p=2^3p+2
cancelling the bases ie 2 we have
(-3)*-2p=3p+2
6p=3p+2
taking like terms we have
6p-3p=2
3p=2
p=2/3